If the equation x^4+ax^2+bx+c=0 has roots 1,2,3 Find C
Answer by stanbon(75887)

If the equation x^4+ax^2+bx+c=0 has roots 1,2,3 Find C
-----------
If x = 1: 1+a + b + c = 0
---
If x = 2: 16 +4a + 2b + c = 0
----
If x = 3: 81 + 9a + 3b + c = 0
------------------------------------
Rearrange and solve for a,b,c
---
a + b + c = 0
4a + 2b + c = -16
9a + 3b + c = -81
---------------------------------
I used a matrix method to get:
a = -(49/2)
b = 115/2
c = -33
===============
Cheers,
Stan H.